OpenStudy (anonymous):

The sum of the squares of 3 consecutive positive integers is 116. What are the numbers? Which of the following equations is used in the process of solving this problem? 3n^2+ 5 = 116 3n^2+ 3n + 3 = 116 3n^2+ 6n + 5 = 116

OpenStudy (blurbendy):

Consecutive integers would be represented by x, x+1, x+2 Squaring and summing them would make x^2+(x+1)^2+(x+2)^2=116 foil and combine like terms.

OpenStudy (raden):

are you sure that is 116 ?

OpenStudy (anonymous):

if what is 116??

OpenStudy (raden):

they are cant be integers

OpenStudy (anonymous):

the question is correct

OpenStudy (raden):

yeah, but the wrong statement. if just findind an equation we can do it, but if want to solving we cant solve it

OpenStudy (anonymous):

I trying to find the process that is used to solve it

OpenStudy (raden):

i meant, can be solved but if just they are real numbers not integer numbrs

OpenStudy (whpalmer4):

None of the equations listed have integer solutions...

OpenStudy (raden):

if just want to find the equation : x^2+(x+1)^2+(x+2)^2=116 x^2 + x^2 + 2x + 1 + x^2 + 4x + 4 =116 3x^2 + 6x + 5 = 116 ( it is C) but if want solve it : 3x^2 + 6x + 5 = 116 3x^2 + 6x + 5 - 116 = 0 3x^2 + 6x - 111 = 0 ------------------ : 3 x^2 + 2x - 37 = 0 this can not be fatored, so obviously the roots not integer numbers

OpenStudy (raden):

in other words, wrong question. heh

OpenStudy (anonymous):

that's what I needed to know thx

OpenStudy (whpalmer4):

One of the shortcomings of learning from a canned course: no opportunity to give the teacher a hard time about a badly written question :-)